zarkam21:

The sum of two consecutive numbers is -37. Find the numbers.

Shadow:

\[n + (n + 1) = -37\]

kittybasil:

Two consecutive numbers means the first number \(x\) plus 1 would equal the second number. Thus, the second number is \(x+1\) The formula would be as follows: Number One + Number Two equals negative 37. Replace that with the variable expressions:\[\huge{\color{royalblue}{x}+\color{crimson}{(x+1)}=-37}\]Now simplify this expression:\[2x+1=-37\]\[2x\cancel{+1\color{olive}{-1}}=-37\color{olive}{-1}\]\[2x=-38\]\[\frac{\cancel{2}x}{\cancel{\color{teal}{2}}}=\frac{-38}{\color{teal}{2}}\]\[\large{x=-19}\]